Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

In coastal towns in the Pacific Rim region, there is a serious risk of flooding by tsunamis (large waves caused by the vibrations of earthquakes or eruptions). Tsunamis travel quickly and are often very destructive. Therefore, warning systems with sirens have been installed so people can evacuate to higher ground in time. When the system is tested, individual sirens sometimes fail. The sirens operate independently of one another (i.e. the failure of one siren does not change the probability that any other siren fails). Suppose that you live on the coast and two sirens can be heard from youa. If the system is activated, and there is a 3% chance of any siren failing, what is the probability that two sirens will fail? (Check your decimals J )

OpenStudy (anonymous):

|dw:1349045303041:dw| Therefore, the probability for either one of them failing should be: .97*.03+.03*.03 = .0891, or 8.91%

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

oh, whoops, I didn't read your question carefully.... It should be just .03*.03 = .00009

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
3 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!